# Discretized Quantum Exhaustive Search for Variational Quantum Algorithms

### v1

@@ 2.4 Efficient Partial Discretized Quantum Exhaustive Search
When choosing $K=\log(n)$, the number of MUB state options becomes $n^{2}+n$, and the number of different qubit splitting choices is (${{n}\choose{\log(n)}}=n^{\log(n)+O(\log(\log(n)))}$. The result is super-polynomial but still sub-exponential.
2.5 MUB States as an Initial Guess for Ansatzes and Optimizers
@@
One possible outcome from the sampling of the cost function can be a good initial guess or a set of such initial guesses to use for the optimization process, which might help avoid barren plateaus and/or local minima and increase the chances of converging into the global minimum. The choice for the state to use as an initial guess can simply be the state that gave the minimum value in the cost function evaluation (as long as no additional prior knowledge is known about the problem or the structure of the optimal state), or the best state among a basis whose states showed large differences between their cost values, hinting for an area within the cost function landscape with high gradients.
@@
To the best of our knowledge, the existing gate-based ansatzes’ circuits do not parameterize all of the quantum gates inside them, and all problem-agnostic ansatzes do not have a set of parameters that set the circuit to be equivalent to the identity. As a result, if a circuit that prepares a desired initial guess is inserted at the start of the ansatz, the first sampling point of the optimization will be different than the desired initial guess, even when the zero parameters vector is used. In order to overcome this, we suggest two methods -
@@
  • Expressive ansatz - Use an expressive ansatz, that can generate (given the relevant parameters vector) each of the MUB states that were used for the energy landscape sampling. Then, when wishing to perform an optimization algorithm and use a MUB state $|\psi\rangle$ as an initial guess, use the parameter vector that makes the ansatz generate $|\psi\rangle$ as the initial guess for the optimization. Examples of using such expressive ansatzes to represent MUB states were presented in section 2.3.
@@
  • shifted MUBs - Choose an ansatz before sampling the energy landscape, let the ansatz be $U(\theta)$. Choose a fixed parameters vector $\theta_{0}$ (for example, the all-zero parameters vector). Applying the unitary operation $U(\theta_{0})$ to a set of MUB states preserves their mutually unbiased properties (states from each basis will remain orthogonal, and the inner product of states from different bases will remain the same). Therefore, instead of sampling the energy landscape using MUB states $|{\psi_{1}}\rangle,|{\psi_{2}}\rangle\dots$, one can use the shifted MUB states $U(\theta_{0})|{\psi_{1}}\rangle,\;U(\theta_{0})|{\psi_{2}}\rangle\dots$ Then, in order to use $U(\theta_{0})|{\psi_{k}}\rangle$ as an initial guess for the optimization, one can use the circuit preparing the MUB state $|{\psi_{k}}\rangle$ followed by the ansatz with the fixed parameters vector $\theta_{0}$.
@@
As many research papers show [18, 45, 28, 38], choosing a good initial guess (or guesses) for the optimization process of a VQA might affect dramatically the convergence of the algorithm, namely, might affect either (or both) the number of iterations required for convergence and the quality of the final result to which the algorithm converges.
@@
For some problems, there are domain-specific initial guesses, which use prior knowledge of the problem. For example, a common initial guess for finding the ground energy of different molecules, which is found using the Variational Quantum Eigensolver, is the Hartree-Fock state (note that finding it is an NP-complete problem). Such initial guesses do not always produce the best results, and are only available for problems in which a prior knowledge of the problem is present.
@@
In this subsection, we suggested a method to search for an initial guess that can be applied to any VQA, without the need for prior knowledge, which is useful for small-size problems. Of course, for larger sizes, prior knowledge or AI/ML tools are highly useful. Then, by using partial MUB states, one can sample a polynomial number of states, in order to find a state which has a better chance to lead to a convergence to an optimal result.
3 Results
@@
We used the method of DQES on various problems. For problems that require up to 3 qubits to encode the cost function, we used all the states of a full MUB set. Calculating the cost function at all of the states of a full MUB set allows us to sample evenly from the entire problem’s landscape. As the number of states in a full MUB set grows exponentially in the number of qubits, for problems that require more than 3 qubits to encode the cost function we used the partial DQES method, by using either 2-qubit MUB states or 3-qubit MUB states and keeping the rest of the qubits of the problem in the zero state. For each MUB state, we have calculated the cost function for this state analytically without the need to measure multiple copies of the state (therefore, the result does not have statistical errors). Such analytical calculation is possible when one has an explicit and efficient representation of the quantum state and the number of non-identity terms in the operators of the cost function is small. When using the partial DQES, for each state we have repeated the calculation for each possible allocation of the MUB qubits as part total qubits number. For example, when using a 3 qubits MUB state for a problem with a size of 4 qubits, we repeated the process 4 times for each 3-qubit MUB state - each time another qubit was kept in the $|0\rangle$ state, while the 3 other qubits were assigned as the MUB state.
@@
For some of the problems, we used the best resulting MUB states as an initial guess for the VQE algorithm. When running the VQE algorithm, having an explicit representation of the quantum state requires exponential (in the number of qubits) memory, therefore calculating the result analytically is possible only for small size problems. The order of the bases in the plots in this section is the same as listed in Chapter 2.3, in Figure 1 (a) and (b). The order of the states inside each basis is by the computational basis states - $|00\rangle,|01\rangle,|10\rangle,|11\rangle$ for the 2 qubits MUB states and $|000\rangle,|001\rangle,|010\rangle,|011\rangle,|100\rangle,|101\rangle,|110\rangle,|111\rangle$ for the 3 qubits MUB states. Each state in the other bases can be defined as the state resulting from evolving one of the computational basis states using the basis transformation matrix (or the quantum circuit representing the basis). The order of the states is defined by the states that are produced from the states in the computational basis in their regular order, transformed to the other basis in this manner.
3.1 Single qubit example
@@
For a single qubit, the canonical MUBs are the $X$, $Y$, and $Z$ bases, and of course these could be transformed via an arbitrary rotation to any other, non-canonical, shifted MUB.
@@
We started with each of the six MUB-states, generated using an ansatz, and we applied the COBYLA optimizer. For each initial state, the first guess is the MUB state, and the next three guesses are provided (with no optimization attempt) by COBYLA to be a modification of a single angle of the three parametrized angles, by a single pre-defined “distance”. Then, starting from iteration number 4, the COBYLA algorithm chooses the next guess. There are two different ways to stop the iterations: either by the optimizer finding no way to improve the current cost value, or by the cost function reaching a pre-defined threshold. We present two examples of reaching the solution — starting from two different MUB states, as well as a closeup on the path in the Hilbert space (in this case, the Bloch sphere) at the last iterations. In figure Fig.2 we see those optimization paths examples.
@@
Figure 2: Optimization paths (energy curves) for 1 qubit search space example. (a) Optimization path obtained by COBYLA for a single-qubit example of the use of MUB states for VQA: the full path, starting from state |ket|$\rangle$| - ⟩. (b) Optimization path obtained by COBYLA for a single-qubit example of the use of MUB states for VQA: the full path, starting from the state $|-_{i}\rangle$. (c) A closeup on reaching the solution for the optimization path presented in (a). (d) A closeup on reaching the solution for the optimization path presented in (b).
3.2 Molecular electronic structure problems
@@
We have calculated the energy of various Hamiltonians, and used a full set of MUB states of 2 qubits for the partial DQES method. In all of these Hamiltonians we used the Born-Oppenheimer approximation and used STO-3G basis as the basis for the orbitals function approximation (solving the FCI problem). The Hamiltonians were converted into spin Hamiltonian using parity transformation [42], and encoded into qubits after utilizing symmetries in the Hamiltonian to reduce the amount of qubits needed in order to encode the solution by 2 (as described in [10, 44]). All of the Hamiltonian coefficient calculations were done using the “qiskit-nature” Python package[40], version 0.4.3. When we have calculated the Hartree-Fock states for comparisons, we used the same parameters (STO-3G basis as basis for the orbitals and the same atomic distance) and used the “pyscf” Python package[43], version 1.6.3, to do the calculation. Some preliminary details are provided in the appendix.
@@
The first Hamiltonian represents the $H_{2}$ molecule at an atomic distance of $0.75[A]$. 2 qubits were required to represent this Hamiltonian, as the relevant orbitals included in the calculation are the 1s orbitals from each $H$ atom, resulting in 4 spin orbitals. By utilizing symmetries we managed to reduce the number of qubits required in the encoding to 2. The resulting Hamiltonian, in the tensored Pauli operators form, is:
$\begin{split}&\mathbb{H}_{H_{2}}^{r=0.75}=-1.05540303\cdot II+0.38874759\cdot IZ-0.38874759\cdot ZI\\ &-0.01117714\cdot ZZ+0.18177154\cdot XX\\ \end{split}$ (3)
@@
The results are shown in Fig.3. The best MUB state result is the same as the Hartree-Fock state energy. By using the Real Amplitudes Hardware Efficient Ansatz with a single layer, and using the 3 best MUB result as an initial state, we were able to achieve final results within chemical accuracy (within $1.6[mH]$ from the FCI results) in all 3 cases. The ansatz is expressive enough to create all the MUB states, so we used the parameters form of the MUB states representation to initialize the ansatz with these states, as shown in Fig. 1 (c). We achieved these results using a noiseless simulator and the COBYLA optimizer. The convergence results (energy curves — change of energy as a function of the iterations in the optimization process) are shown in Fig.4. All 3 cases converged to approximately the same result.
@@
Figure 3: Exhaustive search of all the 2 qubit MUB states and their energy respect to the $H_{2}$ molecule in 0.75 [A] atomic distance.
@@
Figure 4: Convergence results of the optimization process of a VQE algorithm on the $H_{2}$ molecule in 0.75 [A] atomic distance using a noiseless simulator, the COBYLA optimizer, and different initial states. The initial states are the 3 MUB states that had the lowest energy in the landscaping of the molecule. The right figure is a zoom to the final iterations of the optimization.
@@
The second Hamiltonian represents the $HeH^{+}$ molecule at an atomic distance of $1[A]$. 2 qubits were required to represent this Hamiltonian, as the relevant orbitals included in the calculation are the 1s orbitals from both the $H$ atom and the $He$ atom, resulting in 4 spin orbitals. By utilizing symmetries we managed to reduce the number of qubits required in the encoding to 2. The resulting Hamiltonian, in the tensored Pauli operators form, is:
$\begin{split}&\mathbb{H}_{HeH^{+}}^{r=1}=-3.04506092\cdot II+0.50258052\cdot IZ+0.11926278\cdot IX\\ &-0.50258052\cdot ZI+0.11926278\cdot XI-0.13894646\cdot ZZ\\ &-0.11926145\cdot ZX+0.11926145\cdot XZ+0.11714671\cdot XX\\ \end{split}$ (4)
@@
The results are shown in Fig.5. The best MUB state result is the same as the Hartree-Fock state energy. By using the By using the Real Amplitudes Hardware Efficient Ansatz with a single layer, and using the 3 best MUB result as an initial state, we were able to achieve final results within chemical accuracy also for this problem. We achieved these results using the same simulator and optimizer. The convergence results are shown in Fig.6. Although all 3 cases converged to an energy within chemical accuracy, there is a slight energy difference between the converged energies. The state that had the lowest energy in the landscaping ($|01\rangle$) converged to the lowest energy in the optimization process.
@@
Figure 5: Exhaustive search of all the 2 qubit MUB states and their energy respect to the $HeH^{+}$ molecule in 1 [A] atomic distance.
@@
Figure 6: The convergence results (energy curves — change of energy as a function of the iterations) of the optimization process of a VQE algorithm on the $HeH^{+}$ molecule in 1 [A] atomic distance using a noiseless simulator, the COBYLA optimizer, and different initial states. The initial states are the 3 MUB states that had the lowest energy in the landscaping of the molecule. The right figure is a zoom to the final iterations of the optimization.
@@
The third Hamiltonian represents the $LiH$ molecule at an atomic distance of $1.5[A]$. 10 qubits were required to represent this Hamiltonian when taking into consideration all of the relevant orbitals (6 orbitals, which represent 12 spin orbitals). We used the partial DQES method, defined in 2.4. We used 2 qubits MUB states, and kept the rest of the qubits at the zero state. For each experiment, we tested all the possible pairs of qubits to be defined as the qubits of the MUB state. We tested how the results vary when reducing the amount of orbitals which we took into consideration in the encoding. Each removed orbital reduces the amount of needed qubits in the encoding by 2. The FCI result shown in the figures resembles the FCI result of the tested subspace, meaning - when not considering the removed orbitals. The FCI energy result is different when considering more orbitals and different basis functions, as shown in [4].
@@
The results when using 10, 8, 6, 4 and 2 qubits in the encoding (using orbitals [0,1,2,3,4,5], [0,1,2,3,5], [0,1,2,5], [1,2,5], [1,2] respectively) are shown in figure Fig.7. The best MUB state result is the same as the Hartree-Fock state energy in all of those cases.
@@
Figure 7: Exhaustive search of all the 2 qubit MUB states and their energy respect to the $LiH$ molecule in 1.5 [A] atomic distance when considering different number of orbitals in the calculation. (a) considering 2 orbitals in the calculation. 2 qubits are required in this encoding. (b) considering 3 orbitals in the calculation. 4 qubits are required in this encoding. (c) considering 4 orbitals in the calculation. 6 qubits are required in this encoding. (d) considering 5 orbitals in the calculation. 8 qubits are required in this encoding. (e) considering 6 orbitals in the calculation. 10 qubits are required in this encoding.
@@
From these results, it can be seen that the orbitals whose occupation is the most relevant for the optimal result are orbitals 1 and 2. In addition, it can be seen that the most relevant information is found in the computational basis. Both of these results are known when using the prior knowledge of quantum chemistry. Such information might be gained also for problems where no such prior information is found, like some cases of a system of linear equations, which can be solved using the VQLS algorithm. By gaining such information, an initial guess can be made for the algorithm, and a better choice of ansatz can be made (for example - choosing an ansatz that does not use phase gates at all, like the Real Amplitudes ansatz). The results of the LiH experiments are summarized in Fig.8.
Figure 8: Best exhaustive search results for different amounts of qubits used in the encoding compared to the Hartree-Fock energy and the subspace FCI results.
3.3 Transverse-Field Ising Problem

### v2

@@ II.4 Efficient Partial Discretized Quantum Exhaustive Search
When choosing $K=\log(n)$, the number of MUB state options becomes $n^{2}+n$, and the number of different qubit splitting choices is ${{n}\choose{\log(n)}}=n^{\log(n)+O(\log(\log(n)))}$. The result is super-polynomial but still sub-exponential.
Choosing a state that is in one of the mutually unbiased bases chosen as the trivial state for the nKn-K qubits that are in a tensor product with the $K$ qubits encoded in a MUB state, will result in a duplication of states for different qubit allocations. To avoid such duplication, one can choose a trivial state that is not included in the set of states spanning the chosen unbiased bases. One can use a-priori knowledge of the problem to choose this static state or to use a random state instead, which can be a different state for each qubit allocation or each basis. One can also calculate the energy for multiple different static or random states for each MUB state and each qubit allocation. Calculating the energy for each qubit allocation and MUB state for a constant number of different states instead of just one does not change the complexity of the calculation, therefore the calculation will still be sub-exponential.
II.5 MUB States as an Initial Guess for Ansatzes and Optimizers
@@
One possible outcome from the sampling of the cost function can be a good initial guess or a set of such initial guesses to use for the optimization process. These guesses might help avoid barren plateaus and/or local minima and increase the chances of converging into the global minimum. The choice for the state to use as an initial guess can simply be the state that gave the minimum value in the cost function evaluation (as long as no additional prior knowledge is known about the problem or the structure of the optimal state), or the best state among a basis whose states showed large differences between their cost values, hinting for an area within the cost function landscape with high gradients.
@@
To the best of our knowledge, the existing gate-based ansatzes’ circuits (more details about gate-based ansatzes in appendix B) do not parameterize all the quantum gates inside them, and all problem-agnostic ansatzes do not have a set of parameters that set the circuit to be equivalent to the identity. As a result, if a circuit that prepares a desired initial guess is inserted at the start of the ansatz, the first sampling point of the optimization will be different from the desired initial guess, even when the zero parameters vector is used. In order to overcome this, we suggest two methods:
@@
  • Expressive ansatz - Use an expressive ansatz that can generate (given the relevant parameters vector) each of the MUB states that were used for the energy landscape sampling. Then, when wishing to perform an optimization algorithm and use a MUB state $|\psi\rangle$ as an initial guess, use the parameter vector that makes the ansatz generate $|\psi\rangle$ as the initial guess for the optimization Examples of using such expressive ansatzes to represent MUB states were presented in section II.3.
@@
  • Shifted MUBs - Choose an ansatz before sampling the energy landscape, let the ansatz be $U(\vec{\theta})$. Choose a fixed parameters vector $\vec{\theta}_{0}$ (for example, the all-zero parameters vector). Applying the unitary operation $U(\vec{\theta}_{0})$ to a set of MUB states preserves their mutually unbiased properties (states from each basis will remain orthogonal, and the inner product of states from different bases will remain the same). Therefore, instead of sampling the energy landscape using MUB states $|{\psi_{1}}\rangle,|{\psi_{2}}\rangle,\ldots$, one can use the shifted MUB states $U(\vec{\theta}_{0})|{\psi_{1}}\rangle,\;U(\vec{\theta}_{0})|{\psi_{2}}\rangle,\ldots$. Then, in order to use $U(\vec{\theta}_{0})|{\psi_{k}}\rangle$ as an initial guess for the optimization, one can use the circuit preparing the MUB state $|{\psi_{k}}\rangle$ followed by the ansatz with the fixed parameters vector $\vec{\theta}_{0}$.
@@
As many research papers show [20, 54, 33, 47], choosing a good initial guess (or guesses) for the optimization process of a VQA might dramatically affect the convergence of the algorithm. Namely, it might affect either (or both) the number of iterations required for convergence and the quality of the final result to which the algorithm converges.
@@
For some problems, there are domain-specific initial guesses which use prior knowledge of the problem. For example, a common initial guess for finding the ground state energies of different molecules via VQE is the Hartree-Fock state (note that finding it is an NP-hard problem). Such initial guesses do not always produce the best results, and are only available for problems in which a prior knowledge of the problem is present.
@@
In this subsection, we suggested a method to search for an initial guess that can be applied to any VQA, without the need for prior knowledge, which is useful for small-size problems. For larger problem sizes, by using partial MUB states, one can sample a polynomial number of states in order to find a state which has a better chance to lead to a convergence to an optimal result. Of course, for larger problems sizes, methods based on prior knowledge or AI/ML tools can also be highly useful.
III Results
@@
We used the method of DQES on various problems. For problems that require up to 3 qubits to encode the cost function, we used all the states of a full MUB set. Calculating the cost function at all of the states of a full MUB set allows us to sample evenly from the entire problem’s landscape. As the number of states in a full MUB set grows exponentially in the number of qubits, for problems that require more than 3 qubits to encode the cost function we used the partial DQES method, by using either 2-qubit MUB states or 3-qubit MUB states and keeping the rest of the qubits of the problem in the $|0\rangle$ state. For each MUB state, we have calculated the cost function analytically without the need to measure multiple copies of the state (therefore, the result does not have statistical errors). Such analytical calculation is possible when one has an explicit and efficient representation of the quantum state and the number of non-identity terms in the operators of the cost function is small. When using partial DQES, the set of states we used for the calculation is the combination of each $K$-qubit MUB state with each possible allocation of $K$ qubits in the n-qubit register. For example, when using a 3 qubits MUB state for a problem with a size of 4 qubits, we repeated the process 4 times per each 3-qubit MUB state: each time, another qubit was kept in the $|0\rangle$ state, while the 3 other qubits were assigned as the MUB state.
@@
The order of the bases in the plots in this section is the same as listed in Chapter II.3, in Figure 1 (a) and (b). The order of the states inside each basis is by the computational basis states - $|00\rangle$, $\rangle, |10|10\rangle$, $|11\rangle$ for the 2 qubits MUB states and $|000\rangle$, $\rangle, |010|010\rangle$, $1\rangle, |100|1$$0\rangle$, $\rangle, |110|110\rangle$, $|111\rangle$ for the 3 qubits MUB states. Each state in the other bases can be defined as the state resulting from evolving one of the computational basis states using the basis transformation matrix (or the quantum circuit representing the basis). The order of the states is defined by the states that are produced from the states in the computational basis in their regular order, transformed to the other basis in this manner.
III.1 Single qubit example
@@
For a single qubit, the canonical MUBs are the $X$, $Y$, and $Z$ bases. Of course, these could be transformed via an arbitrary rotation to any other, non-canonical, shifted MUB.
@@
We started with each of the six MUB-states, generated using an ansatz, and we applied the COBYLA optimizer. For each initial state, the first guess is the MUB state, and the next three guesses are provided (with no optimization attempt) by COBYLA to be a modification of a single angle of the three parametrized angles, by a single pre-defined “distance”. Then, starting from iteration number 4, the COBYLA algorithm chooses the next guess. There are two different ways to stop the iterations: either by the optimizer finding no way to improve the current cost value, or by the cost function reaching a pre-defined threshold. In our examples, the former is used. We present two examples of reaching the solution — starting from two different MUB states, as well as a closeup on the path in the Hilbert space (in this case, the Bloch sphere) at the last iterations. In figure Fig. 2 we see those optimization paths examples.
@@
Figure 2: Optimization paths (energy curves) for 1 qubit search space example. (a) Optimization path obtained by COBYLA for a single-qubit example of the use of MUB states for VQA: the full path, starting from state $|$-$\rangle$. (b) Optimization path obtained by COBYLA for a single-qubit example of the use of MUB states for VQA: the full path, starting from the state $|-_{i}\rangle$. (c) A closeup on reaching the solution for the optimization path presented in (a). (d) A closeup on reaching the solution for the optimization path presented in (b).
III.2 Molecular electronic structure problems
@@
We have calculated the energy of various Hamiltonians, and used a full set of MUB states of 2 qubits for the partial DQES method. In all of these Hamiltonians we used the Born-Oppenheimer approximation and used the “qiskit-nature” Python package [49] default basis for the orbitals function approximation (STO-3G basis), solving the full configuration interaction (FCI) problem (for more details about the problem definition, see appendix D). For choosing the amount of orbitals that are considered in the calculation (which varies for the different molecules tested), we also used the package’s default for this basis. We used the default “method type” for the driver class of the “qiskit-nature” package, which is a restricted method, that assumes that the energy of the spin up and spin down spin-orbitals of the same orbital is equal, resulting in these spin-orbitals being occupied or empty together. The Hamiltonians were converted into spin Hamiltonians using parity transformation [51], and encoded into qubits after utilizing symmetries in the Hamiltonian to reduce the amount of qubits needed in order to encode the solution by 2 (as described in [12, 53]). All Hamiltonian coefficient calculations were done using the “qiskit-nature” Python package [49], version 0.4.3. When we have calculated the Hartree-Fock states for comparisons, we used the same parameters (STO-3G basis as the basis for the orbitals and the same atomic distance) and used the “pyscf” Python package [52], version 1.6.3, to perform the calculation. Some preliminary details are provided in appendix D.
In the results presented below, we chose Hamiltonians representing the molecules $H_{2}$, $HeH^{+}$ and $LiH$. Given the basis we used (STO-3G) and the “method type” (RHF), the software automatically chooses $k$ lowest-energy orbitals (kk=2,2,6, for the three molecules respectively) meaning $2k$ spin orbitals. This means orbitals [0,1] for $H_{2}$ and $HeH^{+}$ and orbitals [0,1,2,3,4,5] for $LiH$. For $H_{2}$ and $HeH^{+}$ there are just two electrons and both are analyzed, in the four spin-orbitals. For $LiH$, four electrons participated in the calculation.
@@
The first Hamiltonian represents the $H_{2}$ molecule at an atomic distance of $0.75[\text{\AA }]$. The relevant orbitals included in the calculation are the molecular orbitals formed from the 1s orbitals from each $H$ atom, resulting in 4 spin orbitals. Using the parity transformation and utilizing symmetries to reduce the number of qubits required in the encoding, we converted this Hamiltonian into a qubit Hamiltonian of 2 qubits, in the tensored Pauli operators form:
[…]
@@
The results are shown in Fig. 3. The best MUB state result is the same as the Hartree-Fock state energy. By using the Real Amplitudes Hardware Efficient Ansatz with a single layer, and using the 3 best MUB result as an initial state, we were able to achieve final results within chemical accuracy (within $1.6[mH]$ from the FCI results) for all three states. The ansatz is expressive enough to create all of the MUB states, so we used the parameters form of the MUB states representation to initialize the ansatz with these states, as shown in Fig. 1 (c). We achieved these results using a noiseless simulator and the COBYLA optimizer. The convergence results (energy curves) are shown in Fig. 4. All three states converged to approximately the same result.
@@
Figure 3: Exhaustive search of all the 2 qubit MUB states and their energy with respect to the $H_{2}$ molecule in $0.75[\text{\AA }]$ atomic distance.
@@
Figure 4: Convergence results of the optimization process of a VQE algorithm on the $H_{2}$ molecule in $0.75[\text{\AA }]$ atomic distance using a noiseless simulator, the COBYLA optimizer, and different initial states. The initial states are the 3 MUB states that had the lowest energy in the landscape calculation of the molecule. The right figure is a zoom to the final iterations of the optimization.
@@
The second Hamiltonian represents the $HeH^{+}$ molecule at an atomic distance of $1[\text{\AA }]$. The relevant orbitals included in the calculation are the molecular orbitals formed from the 1s orbitals from both the $H$ atom and the $He$ atom, resulting in 4 spin orbitals. By utilizing symmetries we managed to reduce the number of qubits required in the encoding to 2. The resulting Hamiltonian, in the tensored Pauli operators form, is:
[…]
@@
The results are shown in Fig. 5. The best MUB state result is the same as the Hartree-Fock state energy. By using the Real Amplitudes Hardware Efficient Ansatz with a single layer, and using the 3 best MUB result as an initial state, we were able to achieve final results within chemical accuracy for this problem as well. We achieved these results using the same simulator and optimizer. The convergence results are shown in Fig. 6. Although all three states converged to an energy within chemical accuracy, there is a slight energy difference between the converged energies. The state that had the lowest energy in the landscape calculation ($|01\rangle$) converged to the lowest energy in the optimization process.
@@
Figure 5: Exhaustive search of all the 2 qubit MUB states and their energy with respect to the $HeH^{+}$ molecule in $1[\text{\AA }]$ atomic distance.
@@
Figure 6: The convergence results (energy curves) of the optimization process of a VQE algorithm on the $HeH^{+}$ molecule’s Hamiltonian in $1[\text{\AA }]$ atomic distance using a noiseless simulator, the COBYLA optimizer, and different initial states. The initial states are the 3 MUB states that had the lowest energy in the landscape calculation of the molecule. The right figure is a zoom to the final iterations of the optimization.
@@
The third Hamiltonian represents the $LiH$ molecule at an atomic distance of $1.5[\text{\AA }]$. 10 qubits were required to represent this Hamiltonian when taking all the relevant orbitals (6 orbitals, which represent 12 spin orbitals) into consideration. We used the partial DQES method, defined in II.4. We used 2-qubit MUB states, and kept the rest of the qubits at the $|0\rangle$ state. For each experiment, we tested all possible pairs of qubits to be defined as the qubits of the MUB state. We tested how the results vary when reducing the amount of orbitals which we took into consideration in the encoding. Each removed orbital reduces the amount of needed qubits in the encoding by 2. The FCI result shown in the figures resembles the exact ground state energy of the tested subspace - the Hamiltonian formed only by the considered orbitals. The ground state energy changes when considering different orbitals or different basis functions, as shown in [4].
When analyzing the 12 spin-orbitals case, four electrons participated in the calculation. To analyze the options with less orbitals, we used the ActiveSpaceTransformer(num_electrons, num_spatial_orbitals, active_orbitals) class of the “qiskit-nature” package. If one defines num_electrons = 2 without reducing the number of orbitals and not specifying specific active orbitals (active_orbitals = none), only two electrons are analyzed while the other two fill in the lowest level [0], which leads to analysis of orbitals [1,2,3,4,5]. If one define num_orbitals = 2 while also reducing the number of active electrons via num_electrons = 2, the two active electrons are analyzed in orbitals [1,2].
@@
We checked the option with no reduced orbitals, and the option mentioned above, as well as options with specifically chosen orbitals. Thus, the analysis we did was for orbitals [0,1,2,3,4,5], [0,1,2,3,5], [0,1,2,5], [1,2,5] and [1,2]. This leads to using 10, 8, 6, 4, and 2 qubits, respectively, where in each of those, we saved two qubits relative to a naive parity encoding using symmetries [12, 53]. Results are shown in Fig. 7. The best MUB state result is the same as the Hartree-Fock state energy in all of those cases.
@@
Figure 7: Exhaustive search of all the 2 qubit MUB states and their energy with respect to the $LiH$ molecule in $1.5[\text{\AA }]$ atomic distance when considering different number of orbitals in the calculation. The points from different qubit allocations are showed one after the other, while the ordering of the points for allocation is the same as described as the beginning of section III. (a) considering 2 orbitals in the calculation. 2 qubits are required in this encoding. Total of 20 points were calculated. (b) considering 3 orbitals in the calculation. 4 qubits are required in this encoding. 6 qubit allocations were tested, with a total of 120 points calculated. (c) considering 4 orbitals in the calculation. 6 qubits are required in this encoding. 15 qubit allocations were tested, with a total of 300 points calculated. (d) considering 5 orbitals in the calculation. 8 qubits are required in this encoding. 28 qubit allocations were tested, with a total of 560 points calculated. (e) considering 6 orbitals in the calculation. 10 qubits are required in this encoding. 48 qubit allocations were tested, with a total of 900 points calculated.
@@
From these results, it can be seen that the orbitals whose occupation is the most relevant for the optimal result are orbitals 1 and 2. In addition, it can be seen that the most relevant information is found in the computational basis. Both of these results are known when using the prior knowledge of quantum chemistry. Such information might be also gained for problems where no such prior information is found, like some cases of systems of linear equations, which can be solved using the VQLS algorithm. By gaining such information, an initial guess can be made for the algorithm, and a better choice of ansatz can be made (for example - choosing an ansatz that does not use phase gates at all, like the Real Amplitudes ansatz).
III.3 Transverse-Field Ising Problem